A conducting rod of length 4.5 m is sliding downward between two vertical rails without friction. The rails have negligible resistance, but the resistance of the rod is 0.8 Ω. The rails are connected to an 23.3-Ω resistor, and the entire apparatus is placed in a 0.1-T uniform magnetic field. The rod is sliding down at the constant velocity of 7.4 m/s. What is the value of the constant current generated in the rod?
Answer #1

induced emf in the rod = e = B*L*v = 0.1*4.5*7.4 = 3.33 V

and given two resistances are in series

then the net resistance of the circuit is Rnet = 0.8+23.3 = 24.1 ohm

current I = V/Rnet = 3.33/24.1 = 0.138 A